OpenStudy (anonymous):

Divide and check. 9x^2 + 3x – 2 / 3x

OpenStudy (saifoo.khan):

\[\frac{9x^2 + 3x – 2}{3x}\] is it like this?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

lols , is it possible to divide ? i got (3x-1)(3x+2)/3x

OpenStudy (mimi_x3):

Saifoo can help xD

OpenStudy (saifoo.khan):

Sure. help him.

OpenStudy (saifoo.khan):

i think jb copied an incomplete problem.

OpenStudy (mimi_x3):

lols , incomplete xD

OpenStudy (anonymous):

I coppied the correct problem

OpenStudy (saifoo.khan):

there is something missing in the end.

OpenStudy (mimi_x3):

but it cant be simplified further . So how can it be divide

OpenStudy (mimi_x3):

yeah true something is missing xD

OpenStudy (anonymous):

double checking now

OpenStudy (anonymous):

that's the correct problem. nothing else on the end

OpenStudy (mimi_x3):

are yu sure ? i think that is missing either (3x-1) or (3x+2) at the end

OpenStudy (mimi_x3):

it is*

OpenStudy (anonymous):

positive

OpenStudy (mimi_x3):

positive ?

OpenStudy (anonymous):

nothing else on the problem. that's it all there is. there are some others similar to this one

OpenStudy (mimi_x3):

LOL, it cant be divided then xD

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

\[\large \frac{9x^2 + 3x – 2}{3x}\] \[\large \frac{9x^2}{3x} + \frac{3x}{3x} – \frac{2}{3x}\] \[\large 3x + 1 – \frac{2}{3x}\] So \[\large \frac{9x^2 + 3x – 2}{3x}=3x + 1 – \frac{2}{3x}\] Note: the quotient here is 3x+1 and the remainder is -2

OpenStudy (anonymous):

The answer is: 3x+1 with a remainder of -2